# Symmetric Preparation of Systems

### v2

@@ Appendix C Tame Fréchet spaces and the Nash–Moser theorem
(C.1) $\|f\|_{n}\leq C_{n}\|f\|^{\prime}_{n+r}\quad\text{and}\quad\|f\|^{\prime}_{n}\leq C^{\prime}_{n}\|f\|_{n+r}$
@@
(C.2) $\|Lf\|_{n}\leq C_{n}\|f\|_{n+r}$
for any $n\geq b$. If (C.2) holds for any $n$ then $L$ is tame of degree $r$.
Example C.4.
A $C^{\infty}$ linear partial differential operator $L$ of order $m$ on $C^{\infty}$ is a tame of degree $m$ since $\|Lf\|_{n}\leq C_{n}\|f\|_{n+m}$ for the $C^{k}$ norms for any $n$. In particular, multiplication with $C^{\infty}$ functions are tame of degree 0. Solutions to invertible linear systems of smooth equations give tame maps of degree 0 by Leibniz’ rule. Similarly, for any $C^{\infty}$ system $M(x)$, we find that the mapping
$(Q(t,x),R(x))\mapsto G(t,x)=Q(t,x)(t\operatorname{Id}_{n}+M(x))+R(x)$
@@
(C.3) $\|P(f)\|_{n}\leq C_{n}(1+\|f\|_{n+r})$
for all $f\in U$ and all $n\geq b$. If (C.3) holds for any $n$ then $P$ is tame of degree $r$. We say that $P$ is a tame map if $P$ is defined on an open set, is continuous and satisfies a tame estimate in a neighborhood of any point.
As before, if (C.3) holds for any $n$ then we say $P$ satisfies a tame estimate of degree $r$.
Theorem C.11.
@@
The following is the Nash–Moser theorem that we shall use.
Theorem C.12.
@@
Example C.13.
It follows from Corollary B.4 that the mapping
$C^{\infty}(\mathbf{R}\times\mathbf{R}^{n})\times C^{\infty}(\mathbf{R}^{n})\ni(Q(t,x),M(x))\mapsto G(t,x)=P(t,M(x))Q(t,x)+R(x)$
with $M(x)=M^{*}(x)\ \forall\,x$ is a $C^{\infty}$ tame surjective map of degree 0 with local $C^{\infty}$ tame right inverses of degree 4.
@@ Acknowledgment.
I would like to thank Gregory Berkolaiko at Texas A&M, who asked me about symmetric preparation theorems like (1.4) which started this project. We then worked together on the formal power series solution in Section 3. I am also indebted to Dmitry Kerner at Ben-Gurion University, who informed us about the nested Artin Approximation, which gives analytic solutions in the two dimensional case from the formal power series solution, see Remark 2.2.

### v3

@@ Appendix B Tame Fréchet spaces and the Nash–Moser theorem
(B.1) $\|f\|_{n}\leq C_{n}\|f\|^{\prime}_{n+r}\quad\text{and}\quad\|f\|^{\prime}_{n}\leq C^{\prime}_{n}\|f\|_{n+r}$
@@
(B.2) $\|Lf\|_{n}\leq C_{n}\|f\|_{n+r}$
for any $n\geq b$. If (B.2) holds for any $n$ then $L$ is tame of degree $r$.
Example B.4.
A $C^{\infty}$ linear partial differential operator $L$ of order $m$ on $C^{\infty}$ is a tame of degree $m$ since $\|Lf\|_{n}\leq C_{n}\|f\|_{n+m}$ for the $C^{k}$ norms for any $n$. In particular, multiplication with $C^{\infty}$ functions are tame of degree 0. Solutions to invertible linear systems of smooth equations give tame maps of degree 0 by Leibniz’ rule. Similarly, for any $m\in\mathbf{N}$ and $C^{\infty}$ system ${\mathbb{M}}(t,x)$, we find that the mapping
$(Q(t,x),\mathbb{R}(t,x))\mapsto G(t,x)=Q(t,x)(t^{m}\operatorname{Id}_{n}+{\mathbb{M}}(t,x))+\mathbb{R}(t,x)$
@@
(B.3) $\|P(f)\|_{n}\leq C_{n}(1+\|f\|_{n+r})$
for all $f\in U$ and all $n\geq b$. If (B.3) holds for any $n$ then $P$ is tame of degree $r$. We say that $P$ is a tame map if $P$ is defined on an open set, is continuous and satisfies a tame estimate in a neighborhood of any point.
As before, if (B.3) holds for any $n$ then we say $P$ satisfies a tame estimate of degree $r$.
Theorem B.11.
@@
The following is the Nash–Moser theorem that we shall use, which is Theorem 1.7.1 in section III of [5].
Theorem B.12.
@@
Remark B.13.
The local inverse $P^{-1}$ of $P$ in Theorem B.12 inherits the properties of the linear inverse VPVP, since $VP$ is used in the construction of $P^{-1}$, see Theorem 1.7.1 in section III of [5]. For example, if VPVP depends $C^{\infty}$ on some parameters then $P^{-1}$ also depends $C^{\infty}$ on these.
Example B.14.
It follows from Proposition 4.2 that if ${\mathbb{M}}(t)$ is a smooth symmetric $N\times N$ system valued function such that ${\mathbb{M}}(0)=0$ and $\partial_{t}^{m}{\mathbb{M}}\equiv 0$ then the mapping of smooth $N\times N$ system valued functions
(B.4) $C_{0}^{\infty}(\mathbf{R})\times C_{0}^{\infty}(\mathbf{R})\ni(Q(t),{\mathbb{M}}(t))\mapsto\\ G(t)=P(t,{\mathbb{M}}(t))Q(t)+R(t)\in C_{0}^{\infty}(\mathbf{R})\times C_{0}^{\infty}(\mathbf{R})$
is a $C^{\infty}$ tame surjective map of degree 0 with a local $C^{\infty}$ tame right inverse of degree $3+m$.
@@ Acknowledgment.
I would like to thank Gregory Berkolaiko at Texas A&M, who asked me about symmetric preparation theorems like (1.5) which started this project. We then worked on the formal power series solution in the first order case, see [4, Section 2].